# LePaRD: A Large-Scale Dataset of Judicial Citations to Precedent

### v2

@@ Abstract
We present the Legal Passage Retrieval Dataset, LePaRD. LePaRD contains millions of examples of U.S. federal judges citing precedent in context. The dataset aims to facilitate work on legal passage retrieval, a challenging practice-oriented legal retrieval and reasoning task. Legal passage retrieval seeks to predict relevant passages from precedential court decisions given the context of a legal argument. We extensively evaluate various approaches on LePaRD, and find that classification-based retrieval appears to work best. Our best models only achieve a recall of 59% when trained on data corresponding to the 10,000 most-cited passages, underscoring the difficulty of legal passage retrieval. By publishing LePaRD, we provide a large-scale and high quality resource to foster further research on legal passage retrieval. We hope that research on this practice-oriented NLP task will help expand access to justice by reducing the burden associated with legal research via computational assistance. Warning: Extracts from judicial opinions may contain offensive language.
@@ 1 Introduction
To help address the high costs of legal research, and the resulting access to justice issues, and to foster more legal NLP research on practice-oriented tasks, we release the Legal Passage Retrieval Dataset LePaRD. LePaRD represents a large set of previously cited U.S. federal precedent, containing millions of argument contexts and the relevant target passage. In this work, we document the construction of LePaRD and describe relevant dataset statistics. We also extensively evaluate various retrieval approaches from the NLP literature (see e.g., Yang et al., 2017; Reimers and Gurevych, 2019; Mahari, 2021; Tay et al., 2022), some of which have been applied to other legal information retrieval tasks (e.g., Ma et al., 2021a; Rosa et al., 2021). Our most accurate method achieves a recall@10 of 59% on the LePaRD test set, indicating that legal passage retrieval is a challenging task that requires new technical approaches. No large-scale resources for legal passage retrieval exists and we address this gap by constructing and releasing LePaRD.
@@ 2 Related Work
A growing body of work investigates legal citation prediction Dadgostari et al. (2021); Huang et al. (2021) or the retrieval of relevant cases given a query Sansone and Sperlí (2022); Ma et al. (2021b). Based on the preceding context from a legal document, the goal in legal citation prediction is to identify the citation that supports the context in question. By contrast, in legal passage retrieval, the aim is to identify a specific passage of precedent rather than a citation to a whole case (which is usually tens or even hundreds of pages long). We believe there are several reasons to focus on legal passage retrieval over legal citation prediction. Legal citation prediction accuracy numbers seem very strong (see e.g., Huang et al., 2021). We attribute these results to the long-tailed distribution of citations and believe that models take shortcuts to determine a topic for a snippet and then return the most cited cases for these topics—whereas legal passage retrieval inherently requires more involved legal reasoning. This also connects to relevance in legal search, i.e., finding the appropriate target Van Opijnen and Santos (2017). We believe legal relevance is more strongly captured by searching for short passages, rather than predicting citations to entire cases, because a case is likely to deal with multiple independent arguments.
@@ 3.2 Dataset Construction
Length of passage text (chars) 306 225 24 179 253 362 18,342
Length of preceding context (chars) 562 216 5 430 516 638 14,062
Training examples per passage 2.22 8.45 1 1 1 2 4,136
Training examples per destination 6.73 12.2 1 1 3 7 1,375
Training examples per source 11.3 51.8 1 2 4 10 14,450
Training examples per source court 9,831 57,236 1 14 295 1,335 865,594
@@
LePaRD contains the preceding context, target passage, destination court, source court, destination decision date, and source decision date for each quotation that could be matched to a passage. Ultimately, we extract and validate 1.8 million unique target passages that have appeared in approximately 4.3 million contexts.
@@ 4 Dataset Statistics
Quotations serve several purposes in legal writing: they may be used for emphasis, refer to case documents and exhibits, introduce information from witness or expert testimony, cite supporting materials like treatises or academic publications, or they may reference precedential court opinions. Across all federal judicial opinions, we identify 15.8 million quotations over five words long and we successfully match 4.3 million of these to precedential passages. While quotations and citations in judicial opinions offer several interesting avenues for legal NLP and legal passage retrieval, we focus on quotations that can be mapped to single sentences from another opinion. Future work could also examine the retrieval of longer passages or move beyond quotations to general citations (many of which are associated with a “pincite” or page number). In this section, we present several summary statistics about LePaRD (see Table 1) and we highlight some key observations.
First, we note that citations in judicial opinions obey a long-tailed distribution, with the top-1% accounting for 18% of all citations and 64% of all passages receiving just 1 citation. This results in an inherent imbalance in the dataset, raising unique challenges for legal precedent retrieval.
Second, the sentence lengths vary substantially and this results in passages and contexts of varying lengths (the longest passage is over 18,000 characters long). This means that many passages and contexts will be truncated by standard text retrieval approaches.
Third, most destination opinions contain several passages (around 7 on average, but occasionally tens or hundreds). This suggests that there are multiple contexts that occur within a single opinion—something that will be familiar to legal practitioners. In our view, this validates the approach of using local context before a quotation rather than searching for more remote context that may be less relevant (for example, many opinions will discuss factors related to jurisdiction or venue early on but these will not come up anywhere else in the opinion).
Fourth, the average source opinion is represented 11 times in our data. While we treat passages from the same source as separate, it appears likely that they would be conceptually linked (since the portions of an opinion that are cited tend to be somewhat novel or unique and it is uncommon, though not impossible, for there to be multiple such passages in the same opinion). Future work could thus explore whether passage retrieval benefits from grouping passages by their source.
@@ 5 Experiments
Legal passage retrieval seeks to identify passages of precedent given a legal context. In total, we have around 1.3 million unique candidate passages that have been quoted at least once. Hence, given a legal context $x_{i}$, the task is to retrieve the relevant cited passage $y_{i}$ from the set of all possible passages $\{y_{1},y_{2},...,y_{n}\}$.
@@ 5.1 Results
BM25 10K 4.94 19.66 11.38 8.85 5.01 19.52 11.37 8.88
20K 4.18 16.48 9.53 7.40 4.13 16.67 9.56 7.39
50K 3.41 13.25 7.66 5.96 3.41 13.53 7.79 6.04
SBERT 10K 3.78 17.89 9.79 7.34 3.8 17.65 9.75 7.35
20K 3.23 14.13 7.92 6.03 3.08 14.2 7.85 5.92
50K 2.33 10.73 5.91 4.45 2.3 10.67 5.88 4.42
fine-tuned SBERT 10K 8.65 43.95 26.31 19.88 8.99 44.79 26.27 19.84
20K 7.84 38.39 20.8 15.47 7.53 38.5 20.72 15.32
50K 5.12 26.07 13.98 10.31 5.0 25.9 13.78 10.12
LEGAL-BERT Classifier 10K 16.38 47.54 30.66 25.42 16.66 47.65 30.75 25.52
20K 12.52 36.61 23.49 19.43 12.8 36.84 23.77 19.71
50K 8.42 24.99 15.92 13.12 8.77 25.22 16.25 13.48
DistilBERT Classifier 10K 19.39 58.24 37.26 30.73 19.67 59.12 37.73 31.09
20K 16.32 51.75 32.45 26.46 16.74 52.16 32.91 26.94
50K 12.11 38.78 24.21 19.69 12.32 39.34 24.57 20.0
@@
Recall results improve even further if we treat legal passage retrieval as a supervised classification task: Rather than seeking to embed a context and target passage close in some representation space, we assign a unique class label to each passage, and then aim to predict that label from the legal context (see e.g., Mahari, 2021; Tay et al., 2022). We experiment with two different models, and observe that the DistilBERT model achieves the best overall performance in all settings. Our best performance in the 10K label setting suggests that the correct target passage would be predicted among the top 10 search results in 6 out of 10 cases.
@@ Appendix B Further Dataset Statistics
Appendix C Removing Citations in Data
U.S. judicial opinions contain numerous citations to precedent and these frequently also appear in the preceding contexts. We investigate whether removing citations increases model performance (they might be considered noise) or whether citations are a weak form of label leakage (it could be the case that the citations to target passages appear in the preceding context). To test this, we run experiments where citations have been removed from the preceding contexts. We find that results are similar to our main results reported in Table 3.
We extract citations from our data using chatGPT 3.5. We apply the following system prompt: You are a helpful legal assistant with lots of knowledge about the citation style of U.S. court opinions. You are given a short passage from a court opinion. Extract strings containing references to other cases exactly as they appear in the provided text. Extract all text including the case name, citation, pincite, and year, as applicable. Return it as a python list of strings. and add text from LePaRD as the user prompt. We do this for 1,500 examples, and manual inspection yields that this has almost perfect precision, although recall could be improved.
We then use these extracted citations as a dataset, and train a RoBERTa model Liu et al. (2019) in a token-classification manner. The objective is to predict $1$ for each token that is part of a citation, and 0 otherwise. The fine-tuned RoBERTa model achieves an F1 of of 75%, with a precision of 72% and recall of 77%. Manual inspection suggests that the "false positives" predicted by the model are often valid citations which were not extracted by chatGPT in the first place.
We then remove all predicted citations from our dataset containing the 10,000 most cited passages and we re-run our experiments from Section 5. We show results in Table 4.
Approach   Development Set
rc@1 rc@10
BM25 5.2 (+0.26) 20.78 (+1.11)
SBERT 3.95 (+0.17) 18.43 (+0.54)
fine-tuned SBERT 10.41 (+1.85) 47.1 (+3.15)
Classification 18.49 (-0.9) 56.63 (-1.61)
Table 4: Results on dataset with removed citations: Recall at 1, 10 results for development set using 10K labels using various baselines. In brackets, difference (in percentage point) to main results reported in Table 3.
We find that results are generally comparable with those reported in Table 3. Performance increases slightly for BM25 and the cross-encoder (SBERT) experiments, indicating that the citations might be a source of noise. However, performance of the classification experiment Tay et al. (2022) decreases slightly, indicating that citations might also be somewhat correlated with the target label. We underscore that all of these effects are small. Nonetheless, we also release a version of the dataset with the removed citations and our citation identification model to enable others to use citation-free contexts.

### v3

@@ Abstract
We present the Legal Passage Retrieval Dataset, LePaRD. LePaRD contains millions of examples of U.S. federal judges citing precedent in context. The dataset aims to facilitate work on legal passage retrieval, a challenging practice-oriented legal retrieval and reasoning task. Legal passage retrieval seeks to predict relevant passages from precedential court decisions given the context of a legal argument. We extensively evaluate various approaches on LePaRD, and find that classification-based retrieval appears to work best. Our best models only achieve a recall of 38% when trained on data corresponding to the 10,000 most-cited passages, underscoring the difficulty of legal passage retrieval. By publishing LePaRD, we provide a large-scale and high quality resource to foster further research on legal passage retrieval. We hope that research on this practice-oriented NLP task will help expand access to justice by reducing the burden associated with legal research via computational assistance. Warning: Extracts from judicial opinions may contain offensive language.
@@ 1 Introduction
To help address the high costs of legal research, and the resulting access to justice issues, and to foster more legal NLP research on practice-oriented tasks, we release the Legal Passage Retrieval Dataset LePaRD. LePaRD represents a large set of previously cited U.S. federal precedent, containing millions of argument contexts and the relevant target passage. In this work, we document the construction of LePaRD and describe relevant dataset statistics. We also extensively evaluate various retrieval approaches from the NLP literature (see e.g., Yang et al., 2017; Reimers and Gurevych, 2019; Mahari, 2021; Tay et al., 2022), some of which have been applied to other legal information retrieval tasks (e.g., Ma et al., 2021a; Rosa et al., 2021). Our most accurate method achieves a recall of 38% on the LePaRD test set, indicating that legal passage retrieval is a challenging task that requires new technical approaches. No large-scale resources for legal passage retrieval exists and we address this gap by constructing and releasing LePaRD.
@@ 2 Related Work
A growing body of work investigates legal citation prediction Dadgostari et al. (2021); Huang et al. (2021) or the retrieval of relevant cases given a query Sansone and Sperlí (2022); Ma et al. (2021b); Joshi et al. (2023). Based on the preceding context from a legal document, the goal in legal citation prediction is to identify the citation that supports the context in question. By contrast, in legal passage retrieval, the aim is to identify a specific passage of precedent rather than a citation to a whole case (which is usually tens or even hundreds of pages long). We believe there are several reasons to focus on legal passage retrieval over legal citation prediction. Legal citation prediction accuracy numbers seem very strong (see e.g., Huang et al., 2021). We attribute these results to the long-tailed distribution of citations and believe that models take shortcuts to determine a topic for a snippet and then return the most cited cases for these topics—whereas legal passage retrieval inherently requires more involved legal reasoning. This also connects to relevance in legal search, i.e., finding the appropriate target Van Opijnen and Santos (2017). We believe legal relevance is more strongly captured by searching for short passages, rather than predicting citations to entire cases, because a case is likely to deal with multiple independent arguments.
@@ 3.2 Dataset Construction
Length of passage text (chars) 275 192 22 167 233 326 13,091
Length of preceding context (chars) 719 647 5 171 454 1,277 14,275
Training examples per passage 3.62 33.10 1 1 1 3 33,439
Training examples per destination 16.20 30.08 1 3 7 18 2,576
Training examples per source 33.00 245.88 1 3 9 27 93,918
Training examples per source court 35,432 216,851 1 39 798 3,737 3,402,091
@@
LePaRD contains the preceding context, target passage, destination court, source court, destination decision date, and source decision date for each quotation that could be matched to a passage. Ultimately, we extract and validate 3.9 million unique target passages that have appeared in approximately 14 million contexts.
@@ 4 Dataset Statistics
Quotations serve several purposes in legal writing: they may be used for emphasis, refer to case documents and exhibits, introduce information from witness or expert testimony, cite supporting materials like treatises or academic publications, or they may reference precedential court opinions. While quotations and citations in judicial opinions offer several interesting avenues for legal NLP and legal passage retrieval, we focus on quotations that can be mapped to single sentences from another opinion. Future work could also examine the retrieval of longer passages or move beyond quotations to general citations (many of which are associated with a “pincite” or page number). In this section, we present several summary statistics about LePaRD (see Table 1) and we highlight some key observations.
First, we note that citations in judicial opinions obey a long-tailed distribution, with the top-1% accounting for 24% of all citations and 15% of all passages receiving just 1 citation. This results in an inherent imbalance in the dataset, raising unique challenges for legal precedent retrieval.
Second, the sentence lengths vary substantially and this results in passages and contexts of varying lengths (the longest passage is over 13,000 characters long). This means that many passages and contexts will be truncated by standard text retrieval approaches.
Third, most destination opinions contain several passages (around 16 on average, but occasionally tens or hundreds). This suggests that there are multiple contexts that occur within a single opinion—something that will be familiar to legal practitioners. In our view, this validates the approach of using local context before a quotation rather than searching for more remote context that may be less relevant (for example, many opinions will discuss factors related to jurisdiction or venue early on but these will not come up anywhere else in the opinion).
Fourth, the average source opinion is represented 33 times in our data. While we treat passages from the same source as separate, it appears likely that they would be conceptually linked (since the portions of an opinion that are cited tend to be somewhat novel or unique and it is uncommon, though not impossible, for there to be multiple such passages in the same opinion). Future work could thus explore whether passage retrieval benefits from grouping passages by their source.
@@ 5 Experiments
Legal passage retrieval seeks to identify passages of precedent given a legal context. In total, we have around 3.9 million unique candidate passages that have been quoted at least once. Hence, given a legal context $x_{i}$, the task is to retrieve the relevant cited passage $y_{i}$ from the set of all possible passages $\{y_{1},y_{2},...,y_{n}\}$.
@@ 5.1 Results
BM25 10K 9.42 26.52 17.21 14.32 9.62 26.78 17.44 14.54
20K 7.86 23.21 14.82 12.22 8.08 23.36 15.0 12.41
50K 6.77 19.5 12.49 10.33 6.84 19.67 12.58 10.4
SBERT 10K 6.03 19.56 12.01 9.69 6.11 19.75 12.14 9.8
20K 4.92 16.39 9.96 7.99 5.11 16.62 10.18 8.2
50K 4.11 12.95 8.01 8.01 4.16 13.04 8.07 6.56
fine-tuned SBERT 10K 19.18 61.97 38.76 31.54 19.42 62.13 38.91 31.7
20K 15.2 51.87 31.58 25.33 15.41 51.81 31.67 25.46
50K 11.39 39.49 23.74 18.91 11.34 39.52 23.75 18.92
LEGAL-BERT Classifier 10K 35.45 79.22 56.86 49.75 35.68 79.5 57.08 49.96
20K 28.62 68.75 47.74 41.13 28.45 68.68 47.63 41.01
50K 19.32 48.28 32.93 28.13 19.24 48.26 32.92 28.12
DistilBERT Classifier 10K 37.77 81.2 59.23 52.22 38.0 81.23 59.37 52.40
20K 33.07 75.21 53.55 46.68 33.05 74.95 53.49 46.68
50K 26.76 65.53 45.25 38.86 26.63 65.73 45.26 38.82
@@
Recall results improve even further if we treat legal passage retrieval as a supervised classification task: Rather than seeking to embed a context and target passage close in some representation space, we assign a unique class label to each passage, and then aim to predict that label from the legal context (see e.g., Mahari, 2021; Tay et al., 2022). We experiment with two different models, and observe that the DistilBERT model achieves the best overall performance in all settings. Our best performance in the 10K label setting suggests that the correct target passage would be predicted among the top 10 search results in 8 out of 10 cases.